SHOW SOLUTION:


1. Compute the resistance of a conductor given a resistivity of 10.4Ωm, length of 4m, and a diameter of 20mm.


2. A length of a gold wire (α = 3.4 x 10^

-3) has a resistance of 10 ohms at 25 degrees Celsius. Calculate resistance if the temperature rises to 90 degrees Celsius.


1
Expert's answer
2022-03-10T11:15:53-0500

1.

"R=\rho\frac{l}{A}=10.4*\frac{4}{3.14*0.01^2}=1.32*10^5\:\Omega"

2.

"R(t_2)=R(t_1)(1+\alpha(t_2-t_1))\\
=10*(1+3.4*10^{-3}*(90-25))=12\:\Omega"